Observe yourself in a day. find out how many Hours you spend in the following activities: House chores answering learning modules,Planting/Gardening, Using Social media,listening to music,watching television,and sleeping. Record your data. construct a probability Distribution, then compute the mean, variance and standards deviation that You made.


1
Expert's answer
2022-02-21T19:00:12-0500

Solution:

The probability Distribution is as follows-



Mean of Probability Distribution-

"E(X)=\sum X.P(X)=4.6596"

Variance of Probability distribution-

"Var=E(X^2)-[E(X)]^2=26.265-(4.6596)^2=26.265-21.71=4.55"

Standard deviation of probability distribution-

"S.D.=\sqrt{Var}=\sqrt{4.55}=2.13"